# Regularization by regular noise: a numerical result

### v1

@@ 5. Optimality
together with the condition that $\nabla b$ is continuous and bounded, moreover $X_{r}^{n}\rightarrow X_{r}$ and $X_{k_{n}(s)}^{n}\rightarrow X_{s}$ a.s. from Theorem˜2.2, dominated convergence theorem shows as $n\rightarrow\infty$, we have
$c^{n}(t)$ $-\int_{0}^{t}\exp\left(\int_{s}^{t}\nabla b(X_{r})dr\right)\nabla b(X_{s})\{ns\}(b(X_{s})+(B_{s}^{H})^{\prime})\mathrm{d}s\rightarrow 0\quad a.s.$ (5.2)
Besides, observe
$\int_{0}^{t}\exp\left(\int_{s}^{t}\nabla b(X_{r})dr\right)\nabla b(X_{s})\{ns\}(b(X_{s})+(B_{s}^{H})^{\prime})\mathrm{d}s$
$\rightarrow\frac{1}{2}\int_{0}^{t}\exp\left(\int_{s}^{t}\nabla b(X_{r})dr\right)\nabla b(X_{s})(b(X_{s})+(B_{s}^{H})^{\prime})\mathrm{d}s\quad a.s.$
since $\{n\cdot\}$ converges to $\frac{1}{2}$ weakly in $L^{2}([0,t])$ as $n\rightarrow\infty$, the above combining with (5.2) completes the proof. ∎
Remark 5.2.
Notice from Theorem˜5.1, we verify that indeed order $n$ reaches the optimal rate of EM (1.9), although for $b\in{\mathcal{C}}^{1}$ instead of assuming $b\in{\mathcal{C}}^{\alpha}$ for $\alpha<1$. Putting Theorem˜2.2 and Theorem˜5.1 together we can evidently say that order $n$ reaches the optimal (upper bound) rate of EM also for $b\in{\mathcal{C}}^{\alpha}$ simply by the embedding ${\mathcal{C}}^{1}\subset{\mathcal{C}}^{\alpha}$. However, concerning the lower bound of (1.9) for $\alpha<1$ we can not say too much so far.
@@ Acknowledgment
We are grateful to Prof. Rongchan Zhu (Beijing Institute of Technology) for her valuable suggestions.

### v2

@@ 5. Optimality
together with the condition that $\nabla b_{m}$ is continuous and bounded, moreover $\hat{X}_{r}^{n,n}\rightarrow\hat{X}_{r}$ and $\hat{X}_{k_{n}(r)}^{n,n}\rightarrow\hat{X}_{r}$ a.s. from Theorem˜2.2, dominated convergence theorem shows as $n\rightarrow\infty$, we have $^\hat{\mathbb{P}}$-a.s.
$\lim_{n\rightarrow\infty}L_{t}^{m,n}$
$=\lim_{n\rightarrow\infty}n\int_{0}^{t}\int_{0}^{1}\nabla b_{m}(\theta\hat{X}_{r}^{n,n}+(1-\theta)\hat{X}_{k_{n}(r)}^{n,n})\cdot(\hat{X}_{r}^{n,n}-\hat{X}_{k_{n}(r)}^{n,n})\mathrm{d}\theta\mathrm{d}r$
$=\lim_{n\rightarrow\infty}\int_{0}^{t}\int_{0}^{1}\nabla b_{m}(\theta\hat{X}_{r}^{n,n}+(1-\theta)\hat{X}_{k_{n}(r)}^{n,n})\Big(\{nr\}\big(b(\hat{X}_{k_{n}(r)}^{n,n})+(B_{r}^{H})^{\prime}\big)+o(n^{-\varepsilon^{\prime}})\Big)\mathrm{d}r$
$=\frac{1}{2}\int_{0}^{t}\nabla b_{m}(\hat{X}_{r})\big(b(\hat{X}_{r})+(\hat{B}_{r}^{H})^{\prime}\big)\mathrm{d}r=\frac{1}{2}\big(b_{m}(\hat{X}_{t})-b_{m}(x_{0})\big),$ (5.11)
we have applied the fact that $\{n\cdot\}$ converges to $\frac{1}{2}$ weakly in $L^{2}([0,t])$ as $n\rightarrow\infty$ in deriving the penultimate equality and the lase equality holds due to elementary chain rule.
Now we put (5.9), (5.10) and (5) together with taking $m,n\rightarrow\infty$ for both side of (5.8) then get $^\hat{\mathbb{P}}$-a.s.
$\hat{c}(t)=$ $\int_{0}^{t}\hat{c}_{r}\mathrm{d}({\mathcal{H}}b)_{r}^{\hat{X}}+\frac{1}{2}(b(\hat{X}_{t})-b(x_{0})).$
[…]
Meanwhile following from the fact that $\hat{\mathbb{P}}$ a.s., $\tilde{X}^{n}$ converges to $\hat{X}$, $\tilde{c}^{n}$ converges to $\tilde{c}$ and the property of Young’s integral (2.1) we can conclude $\tilde{c}$ also satisfies (2.7) for such $\hat{X}$. It implies that $\hat{c}=\tilde{c}$ a.s. and $^\hat{\mathbb{P}}$-a.s., $\hat{c}^{n}-\tilde{c}^{n}$ converges to 0; hence we get that for given $X$, $c^{n}$ converges to $c$ in probability on $(\Omega,{\mathcal{F}},{\mathbb{P}})$.
Appendix A Auxiliary Proofs
Proof of (4) and (4.21).
Similarly to (4) and (4.6), we have
$\|A_{s,t}\|_{L^{p}_{\omega}}$ $=\|\int_{s}^{t}\mathcal{P}^{H}_{r-[s-(t-s)]}f({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r}^{n})$
$\qquad-\mathcal{P}^{H}_{r-[s-(t-s)]}f({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi^{n}_{k_{n}(r)})\mathrm{d}r\|_{L^{p}_{\omega}}$
$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\sup_{r\in[s,t]}\|\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}\|_{L_{\omega}^{p}}|t-s|^{1-(1-\alpha)H}$
$\leqslant Cn^{-1}\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{1-(1-\alpha)H},$ (A.1)
where in the second inequality we used $\|\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}\|_{L_{\omega}^{p}}\leqslant\|b\|_{{\mathcal{C}}^{0}}n^{-1}$. Hence (4) holds.
${\mathbb{E}}^{s_{1}}\delta A_{s,u,t}=$ ${\mathbb{E}}^{s_{1}}\int_{s_{4}}^{s_{5}}\mathcal{P}_{r-s_{3}}^{H}f({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-\mathcal{P}_{r-s_{3}}^{H}f({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})$
$\qquad\qquad-\mathcal{P}^{H}_{r-s_{3}}f({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n})+\mathcal{P}_{r-s_{3}}^{H}f({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
$+{\mathbb{E}}^{s_{1}}\int_{s_{5}}^{s_{6}}\mathcal{P}_{r-s_{2}}^{H}f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-\mathcal{P}_{r-s_{2}}^{H}f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})$
$\qquad\qquad-\mathcal{P}_{r-s_{2}}^{H}f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r}^{n})+\mathcal{P}_{r-s_{3}}^{H}f({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
$=$ $:I_{1}+I_{2}.$
$\|I_{1}\|_{L_{\omega}^{p}}\leqslant$ $C\|f\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-(1-\alpha)H}\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\big\|_{L_{\omega}^{p}}$
$\qquad+(r-s_{3})^{-(2-\alpha)H}\big\||{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\cdot{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\big\|_{L_{\omega}^{p}}\mathrm{d}r.$ (A.2)
Similar to (4.10), we have
$\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\|_{L^{p}_{\omega}}\leqslant\|{\mathbb{E}}^{s_{1}}|(\varphi_{r}^{n}-\varphi_{r}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{r)}^{n})|\|_{L^{p}_{\omega}}.$ (A.3)
We note that
When $s_{1}\in[k_{n}(r),r],$ we have
[…]
$\leqslant$ $C(r-k_{n}(r)){\mathbb{E}}^{s_{1}}(|B_{k_{n}(r)}^{H}-{\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}|^{\alpha}+|\varphi_{k_{n}(r)}^{n}-\varphi_{s_{1}}^{n}|^{\alpha}).$
Moreover, using (3.5) and (4.2), we have
$\leqslant$ $C(r-k_{n}(r))(|k_{n}(r)-s_{1}|^{\alpha H}+|k_{n}(r)-s_{1}|^{\alpha})\leqslant\frac{C}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|r-s_{1}|^{\alpha}$
where we used the fact $H>1$ in the second inequality. Plugging it into (A.3), we get
$\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\|_{L^{p}_{\omega}}\leqslant\frac{C}{n}|r-s_{1}|^{\alpha}.$ (A.4)
Meanwhile (4.9) and (4.2) yield
${\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}|\leqslant C|r-s_{1}|^{1+\alpha H},$ (A.5)
$\|{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})\|_{L_{\omega}^{p}}\leqslant C\|\varphi_{\cdot}^{n}-\varphi_{k_{n}(\cdot)}^{n}\|_{C_{p}^{0}}\leqslant\frac{C}{n}.$ (A.6)
Applying (A.4), (A.5) and (A.6) into (A.2) gives us
$\|I_{1}\|_{L^{p}_{\omega}}$ $\leqslant C\frac{\|f\|_{{\mathcal{C}}^{\alpha}}}{n}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-(1-\alpha)H}(r-s_{1})^{\alpha}\mathrm{d}r+(r-s_{3})^{-(2-\alpha)H}(r-s_{1})^{1+\alpha H}\mathrm{d}r$
$\leqslant C\frac{\|f\|_{{\mathcal{C}}^{\alpha}}}{n}(t-s)^{(1+\alpha-(1-\alpha)H)\wedge(2+(2\alpha-2)H)}.$ (A.7)
The same bound holds on $I_{2}$. Hence we get (4.21).
@@ Acknowledgment
We are grateful to Prof. Rongchan Zhu (Beijing Institute of Technology) for her fruitful advices. We also are grateful for valuable suggestion on the optimality result from an anonymous reviewer.